### v1

@@ 3.4.4 Decision Margin
$\text{decision\_margin}=p_{f(\mathbf{x}^{\prime})}(\mathbf{x}^{\prime})-t_{f(\mathbf{x}^{\prime})}$ (23)
where $p_{f(\mathbf{x}^{\prime})}(\mathbf{x}^{\prime})$ is the predicted probability and $t_{f(\mathbf{x}^{\prime})}$ is the class-specific threshold (which may differ from 0.5 in multi-label settings).
@@ 4.1 Data
The PTB-XL dataset contains 21 799 clinical 12-lead ECG recordings sampled at 100 Hz as mentioned in [32]. Each recording lasts 10s and includes patient metadata, signal-quality annotations, and diagnostic labels. PTB-XL is available via PhysioNet ([9, 32]) and was first described in Scientific Data by [33].
Table 1 shows the distribution of the five main diagnostic classes. Normal ECGs (NORM) are most common, followed by myocardial infarction (MI), ST/T changes (STTC), conduction disturbances (CD) and hypertrophy (HYP).
@@ 4.2 Data Preprocessing
The authors of PTB-XL provided stratified sample folds, which we adopted without modification. In our case, folds 1–7 was held out for training and fold 9 for validation. For fold 10 we calculated SHAP values, which then were used do create rules in PHAR method and further clustered to obtain medoids (which acted as counterfactuals). Additionally fold 7 was used for calculating statistics for PHAR method and fold 9 was used as a background for calculating SHAP values. Fold 8, not used in any other place, served for queries. This split yielded 15 245 training samples (69.9%), 2 183 validation samples (10.0%), 2 198 samples to create rules and clustering them (10.1%) and 2 173 query samples (10.0%).
@@ 4.3.1 Model architecture
Convolutional blocks extract local waveform features and reduce the temporal dimension by scanning the signal with small filters, stabilizing activations via batch normalization, injecting non-linearity with ReLU, and halving the time length through max-pooling. GRU layer integrates these features over time, retaining important patterns and discarding noise. RealNVP flow warps the embedding distribution into a flexible latent space by alternately scaling and shifting parts of the vector under learned transformations. MLP head maps the latent vector to class logits, which are converted to independent probabilities by sigmoid at inference.
@@ 4.6.1 Rule extraction from SHAP explanations
The following steps follow the PHAR [15] methodology for converting numeric attributions into interval rules.
@@ 4.7.3 MDS Projection and Clustering
For k-means clustering in MDS space, we evaluated $k\in\{2,3,\emph{...},10\}$ using combined metrics of silhouette coefficient and the optimal cluster counts were: NORM (k=k=6), MI (k=k=7), STTC (k=k=6), CD (k=k=5), and HYP (k=k=3). Figure 3 illustrates this process for the CD class, where silhouette analysis revealed k=5 as optimal. The resulting clustering, visualized in Figure 4, shows separation of CD samples into four distinct morphological groups
4.7.4 Medoid Extraction
@@ 5.1.1 Overall Performance Analysis
Table 3 presents the aggregate performance across all 963 test samples, revealing distinct trade-offs between the three approaches. The validity metric, fundamental to any counterfactual explanation, demonstrates perfect performance (1.0) for both Original and Sparse methods, which means generated counterfactuals successfully change the model’s prediction to the desired target class. However, the Aligned Sparse method shows a notable validity drop to 0.8126, representing an 18.74% decrease. This reduction stems from the temporal alignment process, which occasionally produces waveforms that, while physiologically more plausible, fail to cross the model’s decision boundaries for certain classes. For example, when a prototype has fewer beats than the original signal, we prepend zeros to its start to allow a clear overlay on the query sample, but this may negatively affect the counterfactual’s validity.
@@
Stability analysis reveals complementary strengths across methods. Noise stability remains consistently high ($\geq$0.9948) across all approaches, indicating robust performance under measurement noise typical in clinical ECG recordings. Also, temporal stability shows a dramatic improvement from 0.5304 (Original) to 0.7593 (Aligned Sparse), representing a 43.2% enhancement. This improvement is clinically significant, as it indicates that counterfactuals maintain their validity even when subjected to small temporal shifts common in ECG acquisition.
@@ 5.1.2 Class-Specific Performance Analysis
Hypertrophy (HYP) presents the most significant challenge, with validity dropping to only 0.132 in the Aligned Sparse condition. This poor performance reflects several factors: (i) HYP represents the smallest class in our dataset (12.15%), this imbalance leads to limited prototype diversity; (ii) the condition often co-occurs with other pathologies, creating complex multi-label scenarios.
@@ 5.3 Explanation Representation and Expert Feedback
While technical performance metrics provide essential validation of our approach, understanding healthcare professionals’ preferences and cognitive needs is equally critical for successful clinical deployment. The representation and presentation of explanations can significantly impact their utility, regardless of underlying technical quality. To address this challenge, we conducted semi-structured interviews with three practitioners spanning 3 to 40+ years of ECG interpretation experience and varying levels of familiarity with AI systems. Each expert was presented with different visualization modes, including lead-wise importance plots, attention span highlights, and single vs. multiple counterfactual overlays, for identical ECG inputs. They evaluated these modes based on interpretability, cognitive load, clinical relevance, and workflow compatibility.
All experts strongly favored single counterfactual visualizations paired with attention highlighting, citing reduced cognitive effort and improved clarity in identifying diagnostic shifts. Rather than viewing all possible class transitions, they preferred focused, user-driven alternatives relevant to specific clinical scenarios. Experts emphasized the importance of aligning lead-level explanations with domain knowledge and voiced concerns when model attributions diverged from clinically expected leads.
@@ 5.4 Limitations and Challenges
Several limitations emerge from our analysis. The validity-sparsity trade-off in the aligned method represents a fundamental challenge of achieving physiologically realistic counterfactuals while maintaining model prediction changes. This trade-off is particularly pronounced for HYP, suggesting that certain cardiac conditions may be inherently difficult to explain through prototype-based approaches. Class imbalance effects significantly impact performance, with HYP (12.15% of the dataset) showing markedly worse performance than well-represented classes like NORM (43.64%). Future work should investigate targeted sampling strategies and class-aware prototype selection to address these disparities.
The computational complexity of the alignment process, while acceptable for our current implementation, may limit real-time deployment in resource-constrained clinical environments. The DTW-based alignment requires $O(n^{2})$ operations per prototype comparison, potentially limiting scalability to larger prototype databases.
Evaluation scope remains limited to the PTB-XL dataset, and generalization to other ECG datasets, different demographics, or varying acquisition protocols requires further validation. Additionally, our expert evaluation, while promising, requires expansion to larger clinical studies to establish definitive usability and accuracy benchmarks.
@@ 5.5 Future Improvements
The results presented in Section 5.1, demonstrate significant potential for deployment in clinical ECG interpretation systems. The high validity rates for common conditions (MI, CD) and the improved interpretability through sparse modifications address key requirements for AI-assisted diagnosis. The framework’s ability to generate explanations in near real-time makes it suitable for point-of-care applications.
@@
Future improvements should focus on adaptive validity thresholds that balance explanation quality with clinical requirements for different conditions. For highly critical conditions, perfect validity may be essential. Hybrid prototype selection strategies could address class imbalance by combining data-driven clustering with expert-curated examples for rare conditions. This approach would ensure that prototype databases contain representative examples even for uncommon pathologies. And also, real-time adaptation mechanisms could allow the system to learn from clinician feedback, gradually improving prototype selection and counterfactual generation based on expert preferences and clinical outcomes.
The metrics used in our evaluation process capture distinct aspects of counterfactual quality. Similar to [7] and [35], the natural next step is to combine them into a single composite quality metric $Q$ that aggregates four key criteria: validity, sparsity (via the $L_{0}$ ratio), stability, and decision margin. ($\textit{Q}(\mathbf{x}^{\prime})=w_{v}\cdot\text{validity}+w_{s}\cdot(1-\text{sparsity\_ratio})+w_{st}\cdot\text{stability\_score}+w_{m}\cdot\text{decision\_margin}$) The main challenge lies in determining the optimal weights $(w_{v},w_{s},w_{st},w_{m})$, where each weight respectively governs the contribution of each of the aforementioned metrics, for a given clinical application. In future work, we will employ systematic grid search and Bayesian optimization to select weights that best align $Q$ with clinician-assessed explanation quality. Furthermore, integrating a human(experts)-in-the-loop feedback pipeline will allow iterative, live fine-tuning of these weights based on real-world clinical validation.
@@
Our current overlays on ECG plots use SHAP values. To enhance the contrast and readability of ECG feature highlights, we will investigate alternative numeric attribution methods beyond SHAP [13], such as LIME [19], Integrated Gradients [27], DeepLIFT [24], and SmoothGrad [25]. We also plan to explore non-linear transformations of SHAP values or thresholds to enhance contrast and clarity in the visual highlights.
The development of interactive explanation platform will allow clinicians to explore different counterfactual scenarios, adjust sparsity levels, and visualize explanation sensitivity to various modifications. Such platforms could transform counterfactual explanations from static outputs to dynamic diagnostic tools.
@@ 6 Conclusions
This work makes five key contributions. First, we validate counterfactual and heatmap visualizations with domain experts in ECG interpretation. Second, we propose a prototype-based selection method to select representative counterfactual examples of the entire data distribution. Third, we develop alignment and sparsification optimization techniques to map and prune prototypes, yielding clear, query-specific explanations. Fourth, our approach generalizes beyond ECG to other multivariate time-series domains, such as wearable patient vitals, industrial sensor networks, or predictive maintenance systems, since its core components remain model-agnostic and can be adapted with modest domain-specific alignment. Fifth, we adopt evaluation metrics for counterfactual explanations, originally developed in the XAI literature, and tailor them to the unique characteristics of multivariate time-series ECG signals, ensuring meaningful assessment of our method’s fidelity.

### v2

@@ 3.4.4 Decision Margin
$\text{decision\_margin}=p_{\mathcal{M}(\mathbf{x}^{\prime})}(\mathbf{x}^{\prime})-t_{\mathcal{M}(\mathbf{x}^{\prime})}$ (23)
where $p_{\mathcal{M}(\mathbf{x}^{\prime})}(\mathbf{x}^{\prime})$ is the predicted probability and $t_{\mathcal{M}(\mathbf{x}^{\prime})}$ is the class-specific threshold (which may differ from 0.5 in multi-label settings).
@@ 4.1 Data
The PTB-XL dataset contains 21 799 clinical 12-lead ECG recordings sampled at 100 Hz as mentioned in [wagner_2022_ptbxl]. Each recording lasts 10s and includes patient metadata, signal-quality annotations, and diagnostic labels. PTB-XL is available via PhysioNet ([goldberger_2000_physionet, wagner_2022_ptbxl]) and was first described in Scientific Data by [wagner_2020_ptbxl].
Table 1 shows the distribution of the five main diagnostic classes. Normal ECGs (NORM) are most common, followed by myocardial infarction (MI), ST/T changes (STTC), conduction disturbances (CD) and hypertrophy (HYP). Crucially, these diagnostic categories primarily reflect morphological and structural abnormalities (e.g., changes in wave amplitude, segment deviation, or impulse conduction time) rather than chaotic rhythm disorders. Consequently, the vast majority of recordings in this subset retain identifiable cardiac cycles with stable R-peaks, supporting the applicability of the peak-based alignment strategy described in Section 3.
@@ 4.2 Data Preprocessing
The authors of PTB-XL provided stratified sample folds, which we adopted without modification. In our case, folds 1–7 was held out for training and fold 9 for validation. For fold 10 we calculated SHAP values, which then were used do create rules in PHAR method and further clustered to obtain medoids (which acted as counterfactuals). Additionally fold 7 was used for calculating statistics for PHAR method and fold 9 was used as a background for calculating SHAP values. Fold 8, not used in any other place, was held out as an independent test set, providing unseen query samples to evaluate the generalization capability of the entire explanation pipeline. This split yielded 15 245 training samples (69.9%), 2 183 validation samples (10.0%), 2 198 samples to create rules and clustering them (10.1%) and 2 173 query samples (10.0%).
@@ 4.3.1 Model architecture
Convolutional blocks extract local waveform features and reduce the temporal dimension by scanning the signal with small filters, stabilizing activations via batch normalization, injecting non-linearity with ReLU, and halving the time length through max-pooling. GRU layer integrates these features over time, retaining important patterns and discarding noise. RealNVP flow warps the embedding distribution into a flexible latent space by alternately scaling and shifting parts of the vector under learned transformations. While our ablation study indicated comparable classification performance without this block, we retained it to ensure robust modeling of complex probability densities, following the literature discussed in Section 2.3. MLP head maps the latent vector to class logits, which are converted to independent probabilities by sigmoid at inference.
@@ 4.6.1 Rule extraction from SHAP explanations
The following steps follow the PHAR [mozolewski2025phar] methodology for converting numeric attributions into interval rules.
@@ 4.7.3 MDS Projection and Clustering
For k-means clustering in MDS space, we evaluated $k\in\{2,3,\emph{...},10\}$ using combined metrics of silhouette coefficient and the optimal cluster counts were: NORM (k=k=6), MI (k=k=7), STTC (k=k=6), CD (k=k=5), and HYP (k=k=3). Figure 4 illustrates this optimization process for the CD class as a representative case, where the peak at $k=5$ indicates the optimal number of clusters. The resulting clustering, visualized in Figure 5, shows separation of CD samples into four distinct morphological groups
4.7.4 Medoid Selection
@@ 5.1.1 Overall Performance Analysis
Table 3 presents the aggregate performance across all 963 test samples, revealing distinct trade-offs between the three approaches. The validity metric, fundamental to any counterfactual explanation, shows perfect performance (1.0) for both the Original and Sparse methods, indicating that the generated counterfactuals successfully change the model’s prediction to the desired target class. However, the Aligned Sparse method shows a notable decrease in validity to 0.8126, representing an 18.74% drop. This represents an inherent trade-off between improved sparsity and temporal stability, at the cost of reduced validity. Importantly, our framework provides all three variants (Original, Sparse, and Aligned Sparse), allowing users to select the one that best suits their priorities. Perfect validity when explanation correctness is critical, or improved interpretability when physiological plausibility is preferred. For HYP specifically, the severe validity drop (13.2%) suggests the Sparse variant without alignment may be more appropriate until alignment techniques better accommodate this class.
@@
Stability analysis reveals complementary strengths across methods. Noise stability remains consistently high ($\geq$0.9948) across all approaches, indicating robust performance under measurement noise typical in clinical ECG recordings. Also, temporal stability improves dramatically from 0.5304 (Original) to 0.7593 (Aligned Sparse), representing a 43.2% improvement. This improvement is clinically significant, as it indicates that counterfactuals maintain their validity even when subjected to small temporal shifts common in ECG acquisition.
@@ 5.1.2 Class-Specific Performance Analysis
Hypertrophy (HYP), as the minority class in our dataset (12.15%), presents the most significant challenge, with validity dropping to only 0.132 in the Aligned Sparse condition. However, the effects of class representation extend beyond HYP alone. Table 4 reveals a broader pattern: classes with fewer effective prototypes tend to produce lower-validity counterfactuals. After single-label filtering, prototype availability varied substantially, and NORM retained 377 samples, while HYP retained only 23. This affected classes that frequently co-occur with other conditions; for instance, STTC commonly appears alongside HYP (6.92% of recordings, Table 2), reducing the pool of unambiguous examples for both classes. The validity results reflect this pattern: MI achieves 98.9% validity with 160 filtered samples, while STTC (231 samples) and NORM (377 samples) show moderate validity (65.0% and 68.1% respectively), and HYP (23 samples) shows poor validity (13.2%). These findings reveal important limitations in generalizability: prototype-based counterfactual methods require sufficient representation not only in the original dataset but also among single-label instances to capture the morphological diversity within each diagnostic category.
@@ 5.2 Advantages of the prototype-based counterfactuals
Complementing this representative selection, our sparsity optimization ensures that modifications are restricted to the most influential diagnostic regions, preserving the patient’s baseline morphology while allowing the clinician to focus on critical pathological changes.
@@ 5.3 Explanation Representation and Expert Feedback
While technical performance metrics provide essential validation of our approach, it is important to distinguish between two complementary notions of counterfactual quality: (i) validity, which measures whether the counterfactual successfully changes the model’s prediction, and (ii) clinical plausibility, which assesses whether the waveform modifications correspond to physiologically meaningful changes that clinicians would recognize. A counterfactual may achieve perfect validity while producing waveform alterations that lack clinical interpretability, or conversely, suggest clinically meaningful changes that do not cross the model’s decision boundary. To evaluate clinical plausibility and understand healthcare professionals’ cognitive needs, we conducted semi-structured interviews with three practitioners spanning 3 to 40+ years of ECG interpretation experience and varying levels of familiarity with AI systems. Each expert was presented with different visualization modes, including lead-wise importance plots, attention span highlights, and single vs. multiple counterfactual overlays, for identical ECG inputs. They evaluated these modes based on interpretability, cognitive load, clinical relevance, and workflow compatibility. Evaluation protocol is documented in Appendix D.
All experts strongly favored single counterfactual visualizations paired with attention highlighting, citing reduced cognitive effort and improved clarity in identifying diagnostic shifts. Rather than viewing all possible class transitions, they preferred focused, user-driven alternatives relevant to specific clinical scenarios. Regarding clinical plausibility, experts emphasized the importance of aligning lead-level explanations with established pathological mechanisms. For example, when reviewing MI counterfactuals, experts expected modifications concentrated in leads corresponding to affected myocardial regions. They shared concerns when model attributions diverged from these clinically expected patterns, such as when a NORM-to-MI counterfactual modified precordial leads inconsistently with known infarct localization. However, experts also acknowledged cases where model-highlighted regions, while unexpected, revealed subtle waveform features they might have overlooked, suggesting potential for AI-assisted pattern discovery. These observations underscore that validity and clinical plausibility, while related, capture distinct aspects of counterfactual quality.
@@ 5.4 Limitations and Challenges
Several limitations emerge from our analysis. The validity-sparsity trade-off in the aligned method represents a fundamental challenge of achieving physiologically realistic counterfactuals while maintaining model prediction changes. This trade-off reveals important limitations in generalizability that correlate with class representation. While HYP represents the most extreme case (12.15% of the dataset, only 23 samples after filtering, 13.2% validity), the pattern extends to other classes: STTC and NORM achieve only 65.0% and 68.1% validity respectively as target classes, despite NORM being the majority class in the original distribution. This apparent paradox is explained by examining the diversity of prototypes and their co-occurrence patterns. Classes that frequently appear in multi-label combinations (Table 2) lose more samples during single-label filtering, regardless of their original prevalence.
The detailed results in Appendix C (Tables B2–B8) further demonstrate that counterfactual quality degrades in both directions: when generating counterfactuals from minority classes (limited source diversity) and to minority classes (limited prototype options). These findings suggest that prototype-based counterfactual methods have inherent limitations for underrepresented classes in imbalanced, multi-label medical datasets. Future work should investigate targeted sampling strategies, synthetic prototype augmentation, class-aware prototype selection, and relaxed filtering criteria that retain informative multi-label samples to address these disparities.
Furthermore, the reliance on R-peak alignment assumes a definable cardiac cycle, limiting the method’s scope. While effective for the morphological and conduction pathologies (MI, HYP, STTC, CD) examined in this study, this approach faces challenges with complex arrhythmias characterized by irregular rhythms (e.g., atrial fibrillation) or absent R-peaks. Future application to arrhythmia-specific tasks would require integrating rhythm-agnostic alignment strategies.
The computational complexity of the alignment process, while acceptable for our current implementation, may limit real-time deployment in resource-constrained clinical environments. The DTW-based alignment requires $O(n^{2})$ operations per prototype comparison, which can potentially limit scalability to larger prototype databases. Furthermore, our timing measurements (<1<1 second per recording) were obtained on research computing infrastructure and have not been validated on clinical workstations, embedded devices, or higher sampling frequencies (e.g., 500Hz). Scalability testing across diverse deployment environments remains necessary to confirm real-time viability.
Evaluation scope remains limited to the PTB-XL dataset, and generalization to other ECG datasets, different demographics, or varying acquisition protocols requires further validation. Additionally, our expert evaluation involved only three clinicians in semi-structured interviews without a standardized protocol, which limits the generalizability of usability findings. While initial feedback was positive, larger-scale studies with structured evaluation protocols are needed to establish definitive conclusions about clinical utility.
@@ 5.5 Future Improvements
The results presented in Section 5.1, demonstrate significant potential for deployment in clinical ECG interpretation systems. The high validity rates for common conditions (MI, CD) and the improved interpretability through sparse modifications address key requirements for AI-assisted diagnosis. The framework’s ability to generate explanations in near real-time makes it suitable for point-of-care applications, while the use of real-patient medoids facilitates clinical verifiability by referencing actual cases.
@@
Future improvements should focus on adaptive validity thresholds that balance explanation quality with clinical requirements for different conditions. For highly critical conditions, perfect validity may be essential. Hybrid prototype selection strategies could address class imbalance by combining data-driven clustering with expert-curated examples for rare conditions. This approach would ensure that prototype databases contain representative examples, even for uncommon pathologies. In addition, real-time adaptation mechanisms could allow the system to learn from clinician feedback, gradually improving prototype selection and counterfactual generation based on expert preferences and clinical outcomes.
@@
Our current overlays on ECG plots use SHAP values. To enhance the contrast and readability of ECG feature highlights, we will investigate alternative numeric attribution methods beyond SHAP [Lundberg2017], such as LIME [LIME], Integrated Gradients [pmlr-v70-sundararajan17a], DeepLIFT [DeepLIFT_2017], and SmoothGrad [smilkov2017smoothgrad]. We also plan to explore non-linear transformations of SHAP values or thresholds to enhance contrast and clarity in the visual highlights.
The metrics used in our evaluation process capture distinct aspects of counterfactual quality. Similarly to [darias2025evaluating] and [wang2023operationalizing], the natural next step is to combine them into a single composite quality metric $Q$ that aggregates four key criteria: validity, sparsity (through the $L_{0}$ ratio), stability, and decision margin. ($\textit{Q}(\mathbf{x}^{\prime})=w_{v}\cdot\text{validity}+w_{s}\cdot(1-\text{sparsity\_ratio})+w_{st}\cdot\text{stability\_score}+w_{m}\cdot\text{decision\_margin}$) The main challenge lies in determining the optimal weights $(w_{v},w_{s},w_{st},w_{m})$, where each weight respectively governs the contribution of each of the aforementioned metrics, for a given clinical application. In future work, we will employ systematic grid search and Bayesian optimization to select weights that best align $Q$ with clinician-assessed explanation quality. Furthermore, integrating a human(experts)-in-the-loop feedback pipeline will allow iterative, live fine-tuning of these weights based on real-world clinical validation.
The choice of similarity metric for prototype selection may benefit from further investigation. While DTW accommodates heart rate variability, alternative metrics such as correlation-based measures or frequency-domain distances may better capture clinically meaningful similarities for specific diagnostic contexts, and the optimal choice is likely use-case dependent.
Beyond algorithmic metrics, translating these results into clinical practice encounters integration hurdles comparable to those in cybersecurity infrastructure protection [SCHMITT2023100520]. Critical requirements include real-time responsiveness, strict governance, and managing model lifecycle updates amidst changing patient populations. Our framework is designed to meet these demands: sparse, prototype-based explanations are computationally efficient for real-time deployment and provide a transparent audit trail. However, the current implementation represents a foundational algorithmic step. Achieving full clinical utility within a Clinical Decision Support System (CDSS) requires evolving from static outputs to a robust human-in-the-loop workflow.
Future work will therefore prioritize developing an interactive explanation platform that grants clinicians direct control over the decision support process. Such an interface would allow users to explore different counterfactual scenarios, adjust sparsity levels, and visualize the sensitivity of explanations to various modifications. Furthermore, the platform would enable user-controlled multi-label exploration, allowing clinicians to select specific comorbidity combinations as counterfactual targets when clinically relevant, rather than being limited to single-class transitions. This approach addresses the current limitation of fixed prototype selection while maintaining interpretability through user-driven scope control. A systematic evaluation of these features will require extensive user studies to determine the optimal interaction patterns for complex multi-label scenarios.
@@ 6 Conclusions
This work makes five key contributions. First, we validate counterfactual and heatmap visualizations with domain experts in ECG interpretation. Second, we propose a prototype-guided selection method that uses real ECG recordings as counterfactual templates to achieve physiological validity. Third, we develop alignment and sparsification optimization techniques to map and prune prototypes, yielding clear, query-specific explanations. Fourth, our approach generalizes beyond ECG to other multivariate time-series domains, such as wearable patient vitals, industrial sensor networks, or predictive maintenance systems, since its core components remain model-agnostic and can be adapted with modest domain-specific alignment. Fifth, we adopt evaluation metrics for counterfactual explanations, originally developed in the XAI literature, and tailor them to the unique characteristics of multivariate time-series ECG signals, ensuring meaningful assessment of our method’s fidelity.
@@ A.2 Threshold selection
A.3 Ablation Study of the Feature Importance Threshold for PHAR
We performed a sensitivity analysis on a random subset of 350 samples from Fold 10 (the fold dedicated to rule extraction, see Section 4.2) to determine the optimal SHAP feature importance threshold $\tau$ for the Post-hoc Attribution Rule extraction (PHAR) rule generation module. It is important to emphasize that at this intermediate stage of the pipeline, the primary objective is not immediate human interpretability, but rather the preservation of the most critical signal information to define distinct clusters.
Table 6: Impact of the SHAP selection threshold ($\tau$) on the characteristics of generated rules. The study was conducted on a representative subset of $N=350$ samples ($\approx 16\%$ of the test set).
Threshold ($\tau$) Feature Count per Rule ($\mu\pm\sigma$) Inference Time (s)
45th percentile $536.8\pm 458.0$ 359.0
50th percentile $612.1\pm 422.6$ 342.6
70th percentile $634.4\pm 415.2$ 339.5
90th percentile $\mathbf{733.1\pm 284.2}$ 336.7
95th percentile $489.5\pm 186.8$ 332.3
As shown in Table 6 the results highlight two key advantages of the 90th percentile. First, it maximizes pattern granularity: the generated rules contain the highest average number of feature conditions ($733.1$), implying that the algorithm stabilizes a rich set of relevant features rather than discarding them. This high level of detail is crucial for defining homogenous clusters for prototype extraction. Lower thresholds (45–70th percentile) resulted in significantly higher variance in feature counts ($\sigma>415$). This suggests that forcing the model to consider a vast number of low-importance features (background noise) leads to inconsistent feature selection across samples, where the resulting rules fluctuate between broad and narrow definitions.
Second, the 90th percentile improves computational efficiency, reducing the average inference time by approximately 22 seconds per sample compared to the 45th percentile baseline, as the search space for perturbations is more constrained to salient features. In contrast, the 95th percentile led to a sharp decrease in feature count, resulting in oversimplified representations that are less effective for subsequent prototype derivation. Thus, $\tau=90$ was selected to balance descriptive richness with computational performance.